# Diverse Perspectives on AI: Examining People’s Acceptability and Reasoning of Possible AI Use Cases

### v1

@@ D.3. Factors Influencing Moral Foundations in Rationale
Appendix E Survey 2: Explicit Weighing of Harms and Benefits of Use Cases
@@
E.1. Study Overview
To better understand the reasoning behind participants decisions about the judgment and usage of the use cases, we conducted a second study with 1 survey for each category (Labor Replacement Use Cases and Personal Use Cases). The second study includes an additional set of questions to elicit the harms and benefits of developing and not developing an use case to better understand the reasoning behind participants decisions. Furthermore, we asked participants the judgment and usage decision questions before and after the set of harms and benefits questions to see if listing out reasons about an use case would elicit any change in their decisions. The details of the second study can be found in E.2 and the results can be found in E.3.
E.2. Setup and Details
While these same set of questions are asked for all five use cases for our main study, in our second study, participants are randomly allocated a single use case. The second study differs from the main study with an initial set of judgment questions without open-text rationales (Q1 - Initial to Q4 - Initial), which are followed by explicit listing and weighing of the possible harms and benefits of the use case in the context of both developing and not developing the use case. We then again ask participants the same set of judgment questions along with the open-text questions to elaborate on their reasoning, similar to the main study. To understand how the judgment and usage decisions are affected by other factors, we asked the participants about their demographics, ai literacy levels and several other reasoning factors after the main set of questions, and these questions can be found in §A.1 The main questions for the second study can be found in Table 29. The participant demographics for the second study can be found in Tables 30 to 35. The distribution of each use case within each scenario (Labor Replacement Use Cases and Personal Use Cases) for the second study is relatively well-balanced and can be found in Table 28.
@@ E.2. Setup and Details
E.3. Results
To explore the possible impact of explicitly weighing harms and benefits of a use case on participant’s decision, we analyzed the participant’s judgment of acceptability before and after explicit weighing of harms and benefits (Study 2; see § LABEL:sssec:survey-qs for details on questions asked). The Type III ANOVA with Satterthwaite’s method for measurement time (before, after) indicated a marginally significant effect on usage judgment weighed by confidence, which suggests that explicit harms and benefits weighing may have an influence, albeit not at conventional significance levels. We further analyzed reasoning effect on each subset of data pertinent to each use case through a mixed effects regression model with judgment metric as a dependent variable and measurement time as an independent variable with random effect from subject. Interestingly, the result was significant for Customized Lifestyle Coach AI across different judgments including, existence (), confidence-weighed existence (), and confidence-weighed usage judgments (). Explicit weighing also had a significant effect on confidence of existence judgment for Digital Medical Advice AI (). The negative coefficients for Customized Lifestyle AI suggests that weighing harms and benefits caused participants to lower acceptance and positive coefficient to confidence on judgments on Digital Medical AI suggests that weighing harms and benefits solidified decisions. These diverging effects signify an interesting interaction between use cases and explicit weighing of harms and benefits.

### v2

@@ E.3 Factors Influencing Moral Foundations in Rationale
Appendix F Survey 2: Explicit Weighing of Harms and Benefits of Use Cases
@@
F.1 Study Overview
To better understand the reasoning behind participants decisions about the judgment and usage of the use cases, we conducted a second study with 1 survey for each category (Labor Replacement Use Cases and Personal Use Cases). The second study includes an additional set of questions to elicit the harms and benefits of developing and not developing an use case to better understand the reasoning behind participants decisions. Furthermore, we asked participants the judgment and usage decision questions before and after the set of harms and benefits questions to see if listing out reasons about an use case would elicit any change in their decisions. The details of the second study can be found in F.2 and the results can be found in F.3.
F.2 Setup and Details
While these same set of questions are asked for all five use cases for our main study, in our second study, participants are randomly allocated a single use case. The second study differs from the main study with an initial set of judgment questions without open-text rationales (Q1 - Initial to Q4 - Initial), which are followed by explicit listing and weighing of the possible harms and benefits of the use case in the context of both developing and not developing the use case. We then again ask participants the same set of judgment questions along with the open-text questions to elaborate on their reasoning, similar to the main study. To understand how the judgment and usage decisions are affected by other factors, we asked the participants about their demographics, ai literacy levels and several other reasoning factors after the main set of questions, and these questions can be found in §A.1 The main questions for the second study can be found in Table 31. The participant demographics for the second study can be found in Tables 32 to 37. The distribution of each use case within each scenario (Labor Replacement Use Cases and Personal Use Cases) for the second study is relatively well-balanced and can be found in Table 30.
@@ F.2 Setup and Details
F.3 Results
To explore the possible impact of explicitly weighing harms and benefits of a use case on participant’s decision, we analyzed the participant’s judgment of acceptability before and after explicit weighing of harms and benefits (Study 2; see Table31 for details on questions asked). The Type III ANOVA with Satterthwaite’s method for measurement time (before, after) indicated a marginally significant effect on usage judgment weighed by confidence, which suggests that explicit harms and benefits weighing may have an influence, albeit not at conventional significance levels. We further analyzed reasoning effect on each subset of data pertinent to each use case through a mixed effects regression model with judgment metric as a dependent variable and measurement time as an independent variable with random effect from subject. Interestingly, the result was significant for Customized Lifestyle Coach AI across different judgments including, existence (), confidence-weighed existence (), and confidence-weighed usage judgments (). Explicit weighing also had a significant effect on confidence of existence judgment for Digital Medical Advice AI (). The negative coefficients for Customized Lifestyle AI suggests that weighing harms and benefits caused participants to lower acceptance and positive coefficient to confidence on judgments on Digital Medical AI suggests that weighing harms and benefits solidified decisions. These diverging effects signify an interesting interaction between use cases and explicit weighing of harms and benefits.